OpenStudy (anonymous):

Write an equation for the line that passes through (-4,-5) and is perpendicular to y=1/4x-10

OpenStudy (mathstudent55):

What is the slope of the given equation?

OpenStudy (anonymous):

That's all it says

OpenStudy (mathstudent55):

When an equation is written in the form: y = mx + b m is the slope.

OpenStudy (mathstudent55):

In your case, the equation is y = (1/4)x - 10 You have 1/4 multiplying x just like in the form above we have m multiplying x. The slope is m, so the slope of the line is 1/4

OpenStudy (anonymous):

Yeah I got that far. I don't know what equation to write

OpenStudy (mathstudent55):

Ok. The line we are give has a slope of 1/4 The slopes of perpendicular lines are negative reciprocals. That means that the slopes of perpendicular lines multiply to -1.

OpenStudy (mathstudent55):

That also means that if you know the slope of a line and you need the slope of a perpendicular, just take the given slope, write it as a fraction, flip the fraction, and change the sign.

OpenStudy (mathstudent55):

The given line has slope 1/4. That is already a fraction. Now flip the fraction: 4/1 which is the same as 4. Now change the sign: 4 becomes -4.

OpenStudy (mathstudent55):

That means the libe we need has a slope of -4

OpenStudy (mathstudent55):

Now we know the perpendicular will be: y = -4x + b We need to find b.

OpenStudy (mathstudent55):

We now use the point (-4, -5). We replace x and y by -4 and -5, respectively, and solve for b. -5 = -4(-4) + b

OpenStudy (mathstudent55):

-5 = 16 + b -21 = b y = -4x - 21 is the equation we were looking for.

OpenStudy (anonymous):

See now i remember this, thank you so much.

OpenStudy (mathstudent55):

You're welcome.